OpenStudy (anonymous):

solve quadratic equation. t^2=-3t+4

OpenStudy (anonymous):

t^2 + 3t - 4 = 0 (t+4)(t-1)=0 t=-4 or t=1

OpenStudy (nenadmatematika):

correct

jhonyy9 (jhonyy9):

t^2 +3t -4 =0 (t+4)(t-1)=0

OpenStudy (anonymous):

so I set it to zero and go from there. how do I determine if it has any real roots

jhonyy9 (jhonyy9):

the real roots are t+4=0 --- so t=-4 and t-1=0 so t=1

OpenStudy (anonymous):

That is what I though but had to makesure Thanks all